# DRAN: A Distribution and Relation Adaptive Network for Spatio-temporal Forecasting

### v1

@@ IV-B Training Details
After the exploration stage, the averaged hyperparameters—assumed consistent across benchmarks—are presented as follows: The dimension of the adaptive node embedding is 80, consistent with STAEformer [32]. Two-layer MLPs with a hidden dimension of 64 are employed to generate and . In the SFL module, the Conv1d layers are configured with an input channel equal to the length of the lookback window , an output channel of 1, a kernel size of 3, and "circular padding" as defined in the PyTorch package. The Linear layers map the feature dimension to a hidden dimension of 64. For the DSFL module, both the de-stationary attention and spatial attention layers are set to 3. Each de-stationary attention module follows the parameter setup of STAEformer [32], with 4 attention heads and a feed-forward dimension of 256. In the DSFL module, the feature dimensions of and are both set to 160, and the number of attention heads for is also set to 4. In the Stochastic Learner, the latent layers include 3 Linear layers with ReLU activation functions, mapping the features to 64 dimensions. The reconstruction part of the Stochastic Learner consists of 3 Linear layers followed by ReLU activation functions, which remap the feature dimension from 64 to 160. The decoder comprises 2 Linear layers that fuse the deterministic and stochastic representations to produce the target feature outputs. The balance hyper-parameters and are fine-tuned experimentally to account for the stochastic nature and uncertainties of the datasets. Table III presents the variation in prediction error corresponding to different values of , with the value minimizing the error selected as optimal. Moreover, the training process is conducted on the Adam optimizer with a learning rate of 0.001 and a batch size of 32. The number of training epochs is set to 100. We split the datasets into training, validation and test set with a ratio shown in Table I. Numerical experiments for the methods are conducted using various random seeds from the set to obtain the average performance and standard deviation.
@@ IV-C Baselines
We compare our method against several baseline approaches, including state-of-the-art multi-variable time series forecasting methods and spatio-temporal forecasting methods. Spatio-temporal forecasting techniques can be classified into two categories based on how they learn spatial relations: task-adaptive and dynamic-adaptive methods. Task-adaptive methods focus on learning static temporal and spatial relations from training datasets, which remain fixed during the testing phase. In contrast, dynamic-adaptive methods capture dynamically changing relations from input windows or by updating memory with incoming data. Details of the baseline methods are provided in Table II. In multi-variable time series forecasting, DA-RNN is a classic dual-stage attention-based recurrent neural network designed to capture long-term temporal dependencies. InfoTS and AutoTCL focus on enhancing time series representation learning through series augmentations and contrastive learning. InfoTS introduces a novel contrastive learning approach with information-aware augmentations that adaptively select optimal augmentations and a meta-learner network to learn from datasets. AutoTCL achieves unified and meaningful time series augmentations at both the dataset and instance levels, leveraging information theory to enhance representation quality. TGCN, STGCN, DCRNN and GCGRU utilize physical spatial relations as adjacency matrices and employ static neural networks for prediction. TGCN, DCRNN, and GCGRU use Graph Convolutional Networks (GCN) and Recurrent Neural Networks (RNN) to capture spatial and temporal features. STGCN combines temporal and graph convolutions to learn spatial and temporal dependencies. AGCRN utilizes learnable node embeddings to adapt spatial relations to tasks and node-adaptive parameters to capture specific attributes of each node. ASTGCN and STAEformer employ attention mechanisms to capture dynamic changes in input features. RGSL learns spatio-temporal dependencies from a predefined graph and learnable node embeddings. It dynamically fuses features from two graphs using an attention mechanism. Additionally, STAEformer employs learnable node embeddings and concatenates it with input time series to capture static features. MegaCRN utilizes node embeddings to learn the static relations and memory networks to dynamically match sample patterns with learned static features. ST-SSL does not learn static relations and only uses the adjacency matrix based on node distances as a prior graph. It fine-tunes the static graph using node similarities, effectively fusing dynamic and static features. TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph. We evaluate the performance of these models using two metrics: MAE and Mean Absolute Percentage Error (MAPE). The number of samples in the test datasets is denoted as . and denote the predicted and actual observations of spatio-temporal systems.
@@ V-A Comparison Results
Numerical experiments, as shown in Tables IV and V, present the average performance and standard deviation of both the baseline methods and our DRAN model. Our DRAN model outperforms the baseline methods, demonstrating the best overall performance. Specifically, we observe a reduction in MAE of up to 7.8% in weather forecasting and 1.7% in traffic flow prediction compared to the baselines. DRAN achieves the best average performance on the Weather and NYC datasets. In terms of prediction stability, DRAN exhibits slightly larger standard deviations compared to other methods across experiments with different random seeds. This is likely due to the model’s attempt to learn noise components in the representation, which are randomly sampled from distributions. Our method performs less effectively on the MAPE metric but excels on the MAE metric. This suggests that the model prioritizes optimizing overall error rather than minimizing relative error in specific scenarios, such as scenarios with small target values. Consequently, while DRAN’s performance on the MAPE metric is slightly lower than that of other methods, its superior MAE results demonstrate effective overall error control. ST-SSL exhibits strong performance in one-step traffic flow prediction but is less effective on other datasets. This discrepancy may be due to ST-SSL’s emphasis on one-step ahead prediction, as its decoder is not optimized for multi-step forecasting. Additionally, MegaCRN performs well across both tasks, likely due to its task-adaptive node embeddings and the dynamic features learned from its meta memory networks. RGSL benefits from dynamically fusing an explicit prior graph with a learned implicit graph using attention mechanisms. STAEformer performs well on some datasets but poorly on others, which may be attributed to its non-adaptive fusion process. These results suggest that our proposed DRAN model is capable of adapting to various tasks and effectively capturing both static and dynamic features.
@@
| TGCN [28] | MAE | 1.7360.847 | 7.4120.052 | 7.6770.441 | 12.18510.559 | 10.7928.433 | 25.30711.321 | 21.7159.109 |
| MAPE (%) | 0.5920.287 | 34.6110.827 | 35.0851.704 | 37.5328.401 | 36.0457.883 | 44.53610.924 | 45.18110.584 | |
| MemDA [61] | MAE | 1.7860.071 | 6.7770.154 | 7.2460.100 | 6.7800.257 | 6.3440.286 | 20.9391.515 | 20.6311.268 |
| MAPE (%) | 0.6110.026 | 29.3440.543 | 30.3510.126 | 30.0990.342 | 28.5391.064 | 29.6513.243 | 34.9792.950 | |
| ASTGCN [57] | MAE | 1.5210.155 | 6.4810.352 | 6.6980.541 | 8.7235.180 | 7.5483.533 | 15.3895.711 | 12.5003.923 |
| MAPE (%) | 0.5200.054 | 30.7801.221 | 31.2641.976 | 28.6771.634 | 27.1992.117 | 24.7310.916 | 24.4431.578 | |
| TESTAM [60] | MAE | 1.4811.203 | 6.3310.736 | 6.8260.891 | 6.5580.941 | 6.3081.003 | 22.6863.816 | 21.4504.195 |
| MAPE (%) | 0.5070.415 | 30.4182.571 | 31.6282.949 | 30.1603.249 | 29.7054.047 | 32.1435.566 | 37.2377.208 | |
| AGCRN [40] | MAE | 4.8522.311 | 6.3220.847 | 6.5250.510 | 10.8645.159 | 9.8893.836 | 19.7868.580 | 16.4326.402 |
| MAPE (%) | 1.6690.798 | 30.0491.915 | 30.5751.184 | 34.9993.415 | 34.0323.307 | 33.3076.145 | 32.1815.601 | |
| GCGRU [30] | MAE | 1.0120.041 | 5.4570.063 | 5.6310.290 | 7.0363.489 | 6.2502.526 | 12.1532.649 | 10.2771.340 |
| MAPE (%) | 0.3460.015 | 26.8670.700 | 27.2921.476 | 24.3372.990 | 23.7112.231 | 22.3867.043 | 22.8007.607 | |
| DCRNN [58] | MAE | 0.9840.028 | 5.3740.039 | 5.5570.305 | 6.9403.500 | 6.1532.513 | 12.0973.639 | 9.8332.314 |
| MAPE (%) | 0.3360.010 | 26.7720.850 | 27.1651.507 | 24.0703.034 | 23.5652.251 | 21.1823.214 | 21.5983.360 | |
@@
| MegaCRN [59] | MAE | 0.9520.027 | 5.0420.016 | 5.3570.043 | 6.6023.125 | 5.8782.164 | 12.2060.083 | 9.7400.074 |
| MAPE (%) | 0.3250.010 | 25.3290.167 | 26.2750.213 | 23.4233.445 | 22.7102.775 | 18.0310.582 | 17.9720.294 | |
| RGSL [17] | MAE | 0.7270.003 | 5.1490.140 | 5.3350.200 | 6.9213.513 | 6.0822.518 | 13.9451.775 | 11.8592.918 |
| MAPE (%) | 0.2480.001 | 25.6120.183 | 26.1100.924 | 24.1862.957 | 23.2882.249 | 27.14817.848 | 27.37117.905 | |
| ST-SSL [22] | MAE | 1.3940.040 | 5.1350.024 | 5.2650.023 | 5.0420.029 | 4.7140.023 | 12.0100.481 | 9.7900.101 |
| MAPE (%) | 0.4750.013 | 25.4300.295 | 24.6050.265 | 22.6330.112 | 21.8130.808 | 16.3830.100 | 16.8550.228 | |
@@ V-B The Preservation of Spatial Distribution
To evaluate whether the SFL module preserves spatial distributions across various tasks or not, we analyze the spatial distributions of representations before SFL (), after SFL (), and at the forecasting horizon (). The objective is to determine whether the distributions of representations after SFL are closer to those at the forecasting horizon. For each forecasting task, we randomly generate indices and select samples from , , and . We then use Gaussian Kernel Density Estimation to assess the distributions of these representations. As visualized in Fig. 6, the results demonstrate the SFL module’s ability to preserve the spatial distribution of nodes. The primary goal of SFL is to align the spatial distribution of learned representations with that of the forecasting horizon, thereby enhancing accuracy in capturing spatial and temporal distribution changes. In Fig. 6, significant discrepancies between the spatial distribution of representations before SFL and the final horizon are observed. The SFL module effectively reduces these variations in the spatial distribution of learned representations, thereby facilitating the learning process to map spatially preserved representations to predictions.
V-C Dynamic and Static Relations Learning
In Fig. 7, we explore the adaptive process of our model by depicting the dynamic and static adjacency matrices within the DSFL module. Specifically, we showcase these matrices for the first input time step as an example. The dynamic adjacency matrix is derived from the similarity of time series between nodes according to Equation 6, while the static adjacency matrix is obtained by learning the static relations according to Equation 8. Fig. 7 illustrates the spatial relations between nodes of Weather dataset, with darker colors indicating stronger relationships. The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network. The differing concentrations of dynamic and static perspectives allow the model to learn features from various aspects.
@@ V-C Dynamic and Static Relations Learning
| TGCN [28] | MAE | 34.78590.2043 | 33.6049.765 |
| MAPE (%) | 27.9720.754 | 26.1729.541 | |
| GCGRU [30] | MAE | 25.83360.0399 | 22.2658.494 |
| MAPE (%) | 17.7880.442 | 14.3018.571 | |
| STGCN [29] | MAE | 25.30172.3161 | 25.33810.825 |
@@
| TESTAM [60] | 0.075 | 24.303 |
| DRAN (Ours) | 0.075 | 10.737 |
@@ V-D Ablation Studies
The results of the ablation study are depicted in Fig. 9. The final model, incorporating all modules, achieves the best performance. It is evident that the SFL module contributes most significantly to improving prediction accuracy across different tasks. Removing all distribution adaptation modules (w/o Sta) has less impact compared to removing the SFL module (w/o SFL), highlighting SFL as a crucial and indispensable component for spatio-temporal distribution adaptation. The effectiveness of the Stochastic Learner varies depending on the task, aligning with the fact that task uncertainties differ. Comparing the experimental results of "w/o Gate" and "w/o DSFL," the prediction error increases more significantly when the gate mechanism in the DSFL module is removed than when the entire DSFL module is removed. This suggests that fusing dynamic and static representations with a fixed ratio, without considering the time-varying relationships between them, hinders accurate prediction and results in an unsuitable feature combination. Moreover, our model outperforms the scenario where all normalization operations are removed, highlighting the necessity of normalization in spatio-temporal data processing.

### v2

@@ IV-B Training Details
After the exploration stage, the averaged hyperparameters—assumed consistent across benchmarks—are presented as follows: The dimension of the adaptive node embedding is 80, consistent with STAEformer [30]. Two-layer MLPs with a hidden dimension of 64 are employed to generate and . In the SFL module, the Conv1d layers are configured with an input channel equal to the length of the lookback window , an output channel of 1, a kernel size of 3, and "circular padding" as defined in the PyTorch package. The Linear layers map the feature dimension to a hidden dimension of 64. For the DSFL module, both the de-stationary attention and spatial attention layers are set to 3. Each de-stationary attention module follows the parameter setup of STAEformer [30], with 4 attention heads and a feed-forward dimension of 256. In the DSFL module, the feature dimensions of and are both set to 160, and the number of attention heads for is also set to 4. In the Stochastic Learner, the latent layers include 3 Linear layers with ReLU activation functions, mapping the features to 64 dimensions. The reconstruction part of the Stochastic Learner consists of 3 Linear layers followed by ReLU activation functions, which remap the feature dimension from 64 to 160. The decoder comprises 2 Linear layers that fuse the deterministic and stochastic representations to produce the target feature outputs. The balance hyperparameters and are fine-tuned experimentally to account for the stochastic nature and uncertainties of the datasets. Table III presents the variation in prediction error corresponding to different values of , with the value minimizing the error selected as optimal. Moreover, the training process is conducted on the Adam optimizer with a learning rate of 0.001 and a batch size of 32. The number of training epochs is set to 100. We split the datasets into training, validation and test set with a ratio shown in Table I. Numerical experiments for the methods are conducted using various random seeds from the set to obtain the average performance and standard deviation.
@@ IV-C Baselines
We compare our method against several baseline approaches, including state-of-the-art multi-variable time series forecasting methods and spatio-temporal forecasting methods. Spatio-temporal forecasting techniques can be classified into two categories based on how they learn spatial relations: task-adaptive and dynamic-adaptive methods. Task-adaptive methods focus on learning static temporal and spatial relations from training datasets, which remain fixed during the testing phase. In contrast, dynamic-adaptive methods capture dynamically changing relations from input windows or by updating memory with incoming data. Details of the baseline methods are provided in Table II. In multi-variable time series forecasting, DA-RNN is a classic dual-stage attention-based recurrent neural network designed to capture long-term temporal dependencies. InfoTS and AutoTCL focus on enhancing time series representation learning through series augmentations and contrastive learning. InfoTS introduces a novel contrastive learning approach with information-aware augmentations that adaptively select optimal augmentations and a meta-learner network to learn from datasets. AutoTCL achieves unified and meaningful time series augmentations at both the dataset and instance levels, leveraging information theory to enhance representation quality. TGCN, STGCN, DCRNN and GCGRU utilize physical spatial relations as adjacency matrices and employ static neural networks for prediction. TGCN, DCRNN, and GCGRU use Graph Convolutional Networks (GCN) and Recurrent Neural Networks (RNN) to capture spatial and temporal features. STGCN combines temporal and graph convolutions to learn spatial and temporal dependencies. AGCRN utilizes learnable node embeddings to adapt spatial relations to tasks and node-adaptive parameters to capture specific attributes of each node. ASTGCN and STAEformer employ attention mechanisms to capture dynamic changes in input features. RGSL learns spatio-temporal dependencies from a predefined graph and learnable node embeddings, dynamically fusing features from two graphs using an attention mechanism. Additionally, STAEformer employs learnable node embeddings and concatenates it with input time series to capture static features. MegaCRN utilizes node embeddings to learn the static relations and memory networks to dynamically match sample patterns with learned static features. ST-SSL does not learn static relations and only uses the adjacency matrix based on node distances as a prior graph. It fine-tunes the static graph using node similarities, effectively fusing dynamic and static features. TESTAM employs a mixture-of-experts model with three experts: one for temporal modeling, one for spatio-temporal modeling with a static graph, and one for spatio-temporal dependency modeling with a dynamic graph. We also fine-tune the hyperparameters of baseline methods to ensure fairness, especially for datasets not included in their official implementations. We evaluate the performance of these models using two metrics: MAE and Mean Absolute Percentage Error (MAPE). The number of samples in the test datasets is denoted as . and denote the predicted and actual observations of spatio-temporal systems.
@@ V-A Comparison Results
As shown in Tables IV and V, numerical experiments, present the average performance and standard deviation of both the baseline methods and our DRAN model. Our DRAN model outperforms the baseline methods, demonstrating the best overall performance. DRAN achieves the best average performance on the Weather and NYC datasets, and ranks second on the PeMS04 and PeMS08 datasets. Our method performs less effectively on the MAPE metric but excels on the MAE metric. This suggests that the model prioritizes optimizing the overall error rather than minimizing relative error in specific scenarios, such as scenarios with small target values. Consequently, while DRAN’s performance on the MAPE metric is slightly lower than that of other methods, its superior MAE results demonstrate effective overall error control. In terms of prediction stability, DRAN exhibits slightly larger standard deviations compared to other methods across experiments with different random seeds. This is likely due to the model’s attempt to learn noise components in the representation, which are randomly sampled from distributions.
ST-SSL exhibits strong performance in one-step traffic flow prediction but is less effective on other datasets. This discrepancy may be due to ST-SSL’s emphasis on one-step ahead prediction, as its decoder is not optimized for multi-step forecasting. Additionally, MegaCRN performs well across both tasks, likely due to its task-adaptive node embeddings and the dynamic features learned from its meta memory networks. RGSL benefits from dynamically fusing an explicit prior graph with a learned implicit graph using attention mechanisms. STAEformer performs well on some datasets but poorly on others, which may be attributed to its non-adaptive fusion process. These results suggest that our proposed DRAN model is capable of adapting to various tasks and effectively capturing both static and dynamic features.
@@
| TGCN [26] | MAE | 1.7360.847 | 7.4120.052 | 7.6770.441 | 6.1030.106 | 5.7120.072 | 25.30711.321 | 21.7159.109 |
| MAPE (%) | 0.5920.287 | 34.6110.827 | 35.0851.704 | 29.7170.592 | 28.2860.632 | 44.53610.924 | 45.18110.584 | |
| MemDA [76] | MAE | 1.7860.071 | 6.7770.154 | 7.2460.100 | 6.4520.127 | 6.2840.252 | 18.5050.661 | 17.1990.542 |
| MAPE (%) | 0.6110.026 | 29.3440.543 | 30.3510.126 | 29.3120.571 | 27.9950.802 | 28.7191.861 | 32.4631.726 | |
| ASTGCN [72] | MAE | 1.5210.155 | 6.4810.352 | 6.6980.541 | 6.6580.326 | 7.5483.533 | 15.3895.711 | 12.5003.923 |
| MAPE (%) | 0.5200.054 | 30.7801.221 | 31.2641.976 | 28.6741.021 | 27.0281.123 | 24.7310.916 | 24.4431.578 | |
| TESTAM [75] | MAE | 1.4811.203 | 6.3310.736 | 6.8260.891 | 6.4370.852 | 6.2660.901 | 17.6841.814 | 15.9672.758 |
| MAPE (%) | 0.5070.415 | 30.4182.571 | 31.6282.949 | 28.5653.457 | 27.7622.620 | 25.7033.433 | 25.9884.105 | |
| AGCRN [57] | MAE | 1.0940.039 | 5.0660.020 | 5.3660.023 | 5.2790.058 | 4.9410.035 | 12.2630.132 | 9.7200.098 |
| MAPE (%) | 0.3730.013 | 25.5510.200 | 26.5640.209 | 25.4980.268 | 24.6330.286 | 18.2340.324 | 18.2390.278 | |
| GCGRU [28] | MAE | 1.0120.041 | 5.4570.063 | 5.6310.290 | 5.3130.064 | 4.9870.056 | 12.1532.649 | 10.2771.340 |
| MAPE (%) | 0.3460.015 | 26.8670.700 | 27.2921.476 | 25.4280.423 | 24.4900.185 | 22.3867.043 | 22.8007.607 | |
| DCRNN [73] | MAE | 0.9840.028 | 5.3740.039 | 5.5570.305 | 5.3860.036 | 5.0470.031 | 12.0973.639 | 9.8332.314 |
| MAPE (%) | 0.3360.010 | 26.7720.850 | 27.1651.507 | 25.2790.273 | 24.3320.164 | 21.1823.214 | 21.5983.360 | |
@@
| MegaCRN [74] | MAE | 0.9520.027 | 5.0420.016 | 5.3570.043 | 5.1520.049 | 4.8260.021 | 12.0840.158 | 9.5690.175 |
| MAPE (%) | 0.3250.010 | 25.3290.167 | 26.2750.213 | 24.8780.409 | 23.8700.293 | 17.8830.204 | 17.6640.338 | |
| RGSL [16] | MAE | 0.7270.003 | 5.1490.140 | 5.3350.200 | 5.2510.039 | 4.9150.060 | 12.0830.125 | 9.5820.072 |
| MAPE (%) | 0.2480.001 | 25.6120.183 | 26.1100.924 | 24.1862.957 | 23.2882.249 | 18.4300.531 | 18.2050.319 | |
| ST-SSL [20] | MAE | 1.3520.010 | 5.1350.024 | 5.2650.023 | 5.0420.029 | 4.7140.023 | 12.0100.481 | 9.7900.101 |
| MAPE (%) | 0.4610.004 | 25.4300.295 | 24.6050.265 | 22.6330.112 | 21.8130.808 | 16.3830.100 | 16.8550.228 | |
@@ V-B The Preservation of Spatial Distribution
To evaluate whether the SFL module preserves spatial distributions across various tasks or not, we analyze the spatial distributions of representations before SFL (), after SFL (), and at the forecasting horizon (). The objective is to determine whether the distributions of representations after SFL are closer to those at the forecasting horizon. For each forecasting task, we randomly generate indices and select samples from , , and . We then use Gaussian Kernel Density Estimation to assess the distributions of these representations. As visualized in Fig. 6, the results demonstrate the SFL module’s ability to preserve the spatial distribution of nodes. The primary goal of SFL is to align the spatial distribution of learned representations with that of the forecasting horizon, thereby enhancing accuracy in capturing spatial and temporal distribution changes. In Fig. 6, significant discrepancies between the spatial distribution of representations before SFL and the final horizon are observed. The SFL module effectively reduces these variations in the spatial distributions of learned representations, thereby facilitating the learning process to map spatially preserved representations to predictions.
@@
| TGCN [26] | MAE | 34.7850.204 | 24.4480.129 |
| MAPE (%) | 27.9720.754 | 14.5890.219 | |
| GCGRU [28] | MAE | 24.3980.043 | 17.7783.356 |
| MAPE (%) | 15.1050.464 | 10.6222.838 | |
| STGCN [27] | MAE | 25.3022.316 | 25.33810.825 |
@@
| TESTAM [75] | 0.075 | 16.826 |
| DRAN (Ours) | 0.075 | 10.734 |
@@
V-C Dynamic and Static Relations Learning
In Fig. 7, we explore the adaptive process of our model by depicting the dynamic and static adjacency matrices within the DSFL module. Specifically, we showcase these matrices for the first input time step as an example. The dynamic adjacency matrix is derived from the similarity of time series between nodes according to Equation 6, while the static adjacency matrix is obtained by learning the static relations according to Equation 8. Fig. 7 illustrates the spatial relations between nodes of Weather dataset, with darker colors indicating stronger relationships. The dynamic adjacency matrix highlights the strength of relationships between nodes with similar signal patterns, whereas the static adjacency matrix focuses on the signals of individual nodes and some distributed nodes within the network. The contrast in focus enables the model to learn complementary features from both dynamic and static perspectives.
Moreover, the gate mechanism enables adaptive integration of dynamic and static features. We visualize the first channel of the gate signal from three layers of DSFL modules in Fig. 9, illustrating fusion ratio variations on the NYCBike2 test set. Each gate mechanism performs hierarchical fusion, where the blue-filled area represents the contribution of static features to the final deterministic features , and the orange area represents the dynamic features. The fusion ratios vary in sync with the fluctuations of the input inflow and outflow series, demonstrating the gate mechanism’s ability to adaptively balance dynamic and static information.
@@ V-D Ablation Studies
The results of the ablation study are depicted in Fig. 10. The final model, incorporating all modules, achieves the best performance. It is evident that the SFL module contributes most significantly to improving prediction accuracy across different tasks. Removing all distribution adaptation modules (w/o Sta) has less impact compared to removing the SFL module (w/o SFL), highlighting SFL as a crucial and indispensable component for spatio-temporal distribution adaptation. The effectiveness of the Stochastic Learner varies depending on the task, aligning with the fact that task uncertainties differ. Comparing the experimental results of "w/o Gate" and "w/o DSFL," the prediction error increases more significantly when the gate mechanism in the DSFL module is removed than when the entire DSFL module is removed. This suggests that fusing dynamic and static representations with a fixed ratio, without considering the time-varying relationships between them, hinders accurate prediction and results in an unsuitable feature combination.
Moreover, our model outperforms the scenario where all normalization operations are removed, highlighting the necessity of normalization in spatio-temporal data processing.